OpenStudy (campayne):

Use elimination to solve the system of equations. Enter your answer as an ordered pair; example: (10, -25) {2x - 5y = -5 -2x + 8y = -58

OpenStudy (sushi121212):

All you have to do is cancel out each equation: 2x - 5y = -5 -2x + 8y = -58 ___________ 3y = -63 __ ___ 3 3 y= -21 Then just plug it back in: 2x-5(-21)=-5 2x+105=-5 2x=-110 x=-55 Then check: -2x + 8y = -58 -2(-55)+8(-21)=-58 110-168=-58 -58=-58 So your answer is: (-55, -21)

OpenStudy (campayne):

Je vous remercie